# Reverse Question Answering: Can an LLM Write a Question so Hard (or Bad) that it Can’t Answer?

### v1

@@ 6 Limitations
LLMs are sensitive to prompt formats Sclar et al. (2023), so varying prompts could impact LLM accuracy in RQA and QA. To ensure our prompts are reliable, we followed best practices Schulhoff et al. (2024) and kept refining prompts as LLM errors surfaced; the full prompt engineering process is documented in Appendix A.2. Our final prompts will be released and are considered very reasonable implementations of RQA and QA. Further, in Appendix A.6, we test if common prompt engineering strategies (few-shot exemplars, chain-of-thought) can alleviate the low numerical RQA accuracy of GPT-4 but find minimal benefits, suggesting that accuracy gaps between QA and RQA cannot be attributed to prompt formatting alone.
@@ 7 Ethical Considerations
RQA uses abduction, a core reasoning strategy that arrives at a plausible explanation given a set of facts. However, our current findings suggest that LLM abductive reasoning in numerical settings is highly unreliable. We advise practitioners to take caution when using LLMs to perform numerical abductive reasoning in downstream tasks, including designing math exam questions, explaining financial forecasts, proposing economic policies, or diagnosing medical patients from numerical data.
8 Acknowledgements
We would like to thank the CLIP lab at University of Maryland and our external collaborators for their feedback, including Neha Srikanth, Jack Wang, Vishakh Padmakumar, Haozhe An, Yu (Hope) Hou, Abhilasha Sancheti, Seraphina Goldfarb-Tarrant, Aidan Peppin, and Connor Baumler. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber), iis-2339746 (Rudinger), and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research. Access to Command-R was made possible through the Cohere for AI Research Grant Program.
@@ A.2 Prompting Details
Our initial RQA prompt is in Prompt A.8.4. With this prompt, our LLMs generated verbose answers, so we added the instruction that all questions must be “one-sentence” (Prompt A.8.4). Next, we found that it was difficult to reliably parse the question from the model’s generated output, so we added formatting constraints (Prompt A.8.4). At this point, when we looked at the model’s generated questions more closely, we found that models could cheat—adding the answer in the question itself (e.g. giving the question “How many of the 150 people attended the conference” for the answer “150 people”). Thus, we added an instruction to forbid this behavior (Prompt A.8.4). Finally, as we noticed many of the questions were inaccurate, we wanted to study if abstention could alleviate these issues, so we added an instruction (Prompt A.8.4) allowing the model to respond with “IDK” section 2.2. We added abstention to test LLM calibration Feng et al. (2024), but abstention rates are only 3% in QA and <1% in QG, so we do not study it in this work. We keep abstention to avoid re-running all LLMs and omit rare cases of abstention. Our final RQA prompt is shown in Prompt A.8.4.
We then designed our QA prompt by mimicking the format of the final RQA prompt, shown in Prompt A.8.4. We initially wrote the constraint that the answer must be “short” and “just a few words,” but we found these instructions to be ambiguous, and the easy and hard entities split of our dataset had answers that were longer than just a few words; as a result, we removed these instructions, and used “the” instead of “a” to make it clear that there is only one valid answer (Prompt A.8.4). After removing these instructions, we found that models would often generate very long explanations before or after answering the question. To avoid this, we added an instruction stating that we were just looking for the answer and no explanation (Prompt A.8.4). Our final QA prompt is shown in Prompt A.8.4.
@@ A.3 Model Details
LLaMA, Mistral, and Yi models are accessed via huggingface, and all other models are accessed through their respective API endpoints. We allocated 8 NVIDIA:A6000s for Mixtral-8x22B, 8 NVIDIA:A5000s for Mixtral-8x7B, Yi-34B, and LLaMA-70B, 2 NVIDIA:A6000s for Yi-9B and LLaMA-8B, and 1 NVIDIA:A6000 for all other non-API models (which were run on CPU only). Each model was allocated 24 hours to perform both QA and RQA on our dataset.
@@ A.4 Metric Details
To design a metric for QA accuracy, we consider seven answer equivalence metrics, which check if a candidate answer $a_{cand}$ is semantically equivalent to a ground-truth answer $a_{true}$: 1) DSPy-optimized GPT-4o; 2) A rule-based method designed specifically for each dataset; 3) Exact match; 4) Token F1 score; 5) Token Recall Score; 6) Token Precision Score; and 7) PEDANTS Li et al. (2024b), a classifier designed for answer equivalence. The DSPy method in (1) uses a maximum of 10 bootstrapped demos, a maximum of 10 labeled demos, and 20 candidate programs; it uses 64 examples for training (seeding the prompts) and 64 examples for validation. We find optimal decision thresholds for (4), (5), and (6) using the 64 validation examples. We present the agreement with human annotations of each metric in Table 3, which is how we decided which metric to use for each dataset split. Overall, our QA accuracy metric has 94% raw agreement with humans on 1152 held-out examples.
Since there are no automated metrics to check whether a question $q$ can correctly be answered by an entity $a$, we design our own metric for RQA accuracy. Given the strength of the DSPy GPT-4o approach in QA accuracy, we similarly design a DSPy-optimized GPT-4o classifier that determines if $q$ is correctly answered by $a$, using the same hyperparameters for QA accuracy. Overall, this RQA accuracy metric has 90% raw agreement with humans on 1152 held-out examples. We also considered Jury approaches Verga et al. (2024), which ensemble multiple LLMs instead of relying just on a single LLM. However, we found that using majority vote with three/five LLMs boosted our metric’s accuracy by less than 2%, which we did not feel justified the much larger computational expenses.
@@ A.6 RQA with Prompting Engineering
To explore if RQA weaknesses can simply be alleviated with prompt engineering efforts Schulhoff et al. (2024), we test three prompting strategies: 1) Zero-Shot Chain-of-Thought Prompting (asking the LLM to “Think step by step” before answering); 2) Self-Verification (asking the LLM to “Check if the question is accurate after generating a question”); and 3) Five-Shot Prompting (including five exemplars showing the model how to generate a question for an answer). To write exemplars for (3), we pick question/answer pairs when RQA succeeds in the zero-shot setting to try and make the priors in the exemplars most similar to the model’s original generations. The prompts for (1), (2), and (3) are in Prompts A.8.4, A.8.4, and A.8.4, respectively.
We experiment with GPT-4 on Numbers and Numbers+Text, as the model showed a surprising RQA weakness in these settings. GPT-4 is also considered to respond well to prompt engineering efforts, making it a suitable candidate for our prompting strategies. Overall, none of these prompting strategies can close the accuracy gap between RQA and QA (Figure 7). Chain-of-thought prompting increases GPT-4’s RQA accuracy by $\sim 0.15$, but it is still significantly lower than QA, which does not use chain-of-thought. This shows that the accuracy gap between QA and RQA may be a fundamental reasoning flaw of current LLMs that cannot be fully mitigated through prompt engineering.
@@ A.7 Full Consistency Analysis
In Figure 8, we report the full consistency analysis for all 16 of our LLMs and all truth table scenarios. First, we note that the four categories reported in Figure 3 encompass the majority of the truth table. Second, even for smaller LLMs, our claims hold; LLMs can often detect their own question inaccuracies from RQA through QA.
@@ A.8.1 Are RQA questions unanswerable?
We now seek to understand the types of RQA questions generated in the Number+Text setting, complementing our analysis in section 3.4. The Number+Text questions have higher variance and cannot be as neatly categorized as in section 3.4 (e.g. single-step computation). So instead, we study the answerability of 30 generated questions from each LLM, i.e., if the question is clear but leads to an incorrect answer, or if there is an issue with the question that makes it difficult to answer. We adopt five categories of unanswerable questions from Rogers et al. (2023):
@@
2) No Consensus on the Answer: there is not a single, agreed-upon answer to the question. For example, LLaMA generates the question What is the unique property of the Lie algebra E8 that makes it particularly interesting in theoretical physics?, but there are many unique, interesting properties of Lie algebra that would answer the question.
3) Information not yet Discovered: the answer to this question is currently unknown. For example, GPT-4 generates the question How long, in terms of word count, is the sentence that holds the record for being the longest in the English language without using any punctuation?, but it is currently unknown what could theoretically be the longest sentence.
4) Missing Information: There is not enough information in the question, or it is too vague. For example, GPT-4 generates the question How many individuals attended the annual community festival last year according to the final headcount?, which cannot be answered without knowing more details.
@@
As expected, when RQA succeeds, questions are mostly answerable (Figure 9). However, we find that a non-trivial proportion of generated questions when RQA fails are unanswerable, reaching nearly 60% for GPT-4. The most common types of unanswerable questions are those that are missing information, meaning that they are too vague or ambiguous, or those that have false premises or assumptions. While several works explore methods to answer ambiguous questions Min et al. (2020); Kim et al. (2023a) or questions with false presuppositions Yu et al. (2023); Kim et al. (2023b), our analysis reveals a need to avoid generating ambiguous or faulty-presupposition questions in RQA.
@@ A.8.2 Do LLMs give the same RQA questions?
We also note that interestingly, models in the same family will more likely generate duplicated questions. For example, GPT-3.5, GPT-4, and GPT-4o generate the same questions in RQA more often than when compared to other LLM families. Thus, we speculate that these model families likely share similar pre-training and alignment data, which is optimized on through different training recipes.
@@ A.8.4 Are any RQA questions memorized?
Since the presence of duplicates in Appendix A.8.3 suggests that LLMs may just be retrieving similar questions from pretraining rather than reasoning towards new questions in RQA, we now investigate the novelty of the RQA questions Merrill et al. (2024), i.e., whether they are exactly copied from pretraining. We do not know which corpora all of our LLMs are trained on, so we use the Dolma Soldaini et al. (2024) corpus as a proxy for pretraining data. For each generated RQA question $\hat{q}$, we compute how frequently the exact question $\hat{q}$ appears in Dolma via infini-gram Liu et al. (2024b).
Table 7 reveals that in total, $2.87$% of RQA generated questions are exactly found in Dolma. For comparison, $1.25$% of our ground-truth questions are present in Dolma. While we did not explicitly instruct the model to generate a new question that it has not seen in pretraining, practitioners may need to design specialized techniques if they desire more novel questions from RQA.
When comparing exact question match frequency by model, weaker/smaller LLMs tend to copy more from pretraining data, suggesting that smaller LLMs are more prone to RQA memorization. Further, the Hard Fact setting is much less prone to question copying in RQA, likely because the RQA input answers have very low pretraining token count (section 3.3), which also further confirms that LLMs may struggle to retrieve exact pretraining knowledge for long-tail facts Kandpal et al. (2023).

### v2

@@ 6 Limitations
llms are sensitive to prompt formats Sclar et al. (2024b), so varying prompts could impact llm accuracy in RQA and QA. To ensure our prompts are reliable, we followed best practices Schulhoff et al. (2024) and kept refining prompts as llm errors surfaced; the full prompt engineering process is documented in Appendix A.2. Our final prompts will be released and are considered very reasonable implementations of RQA and QA. Further, in Appendix A.6, we test if common prompt engineering strategies (few-shot exemplars, chain-of-thought) can alleviate the low numerical RQA accuracy of GPT-4 but find minimal benefits, suggesting that accuracy gaps between QA and RQA cannot be attributed to prompt formatting alone.
@@ 7 Ethical Considerations
RQA uses abduction, a core reasoning strategy that aims to arrive at a plausible explanation given a set of facts. However, our current findings suggest that llm abductive reasoning in numerical settings is highly unreliable. We advise practitioners to take caution when using llms to reason via numerical abduction in downstream tasks, including designing math exam questions, explaining financial forecasts, proposing economic policies, or diagnosing medical patients from numerical data.
Acknowledgements
We would like to thank the clip lab at University of Maryland and our external collaborators for their feedback, including Neha Srikanth, Haozhe An, Yu Hou, Abhilasha Sancheti, Connor Baumler, Seraphina Goldfarb-Tarrant, Aidan Peppin, Jack Wang, and Vishakh Padmakumar. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber), iis-2339746 (Rudinger), dms-2134012 (Ravichander), and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research. Access to Command-R was made possible with a Cohere for AI Research Grant.
@@ A.2 Prompting Details
Our initial RQA prompt is in Prompt A.8.4. With this prompt, our llms generated verbose answers, so we added the instruction that all questions must be “one-sentence” (Prompt A.8.4). Next, we observed that it was difficult to reliably parse the question from the model’s generated output, so we added formatting constraints (Prompt A.8.4). At this point, when we looked at the model’s generated questions more closely, we saw that models could cheat—adding the answer in the question itself (e.g. giving the question “How many of the 150 people attended the conference” for the answer “150 people”). Thus, we added an instruction to forbid this behavior (Prompt A.8.4). Finally, as we noticed many of the questions were inaccurate, we wanted to study if abstention could alleviate these issues, so we added an instruction (Prompt A.8.4) allowing the model to respond with “IDK” section 2.2. We added abstention to test llm calibration Feng et al. (2024), but abstention rates are only 3% in qa and <1% in RQA, so we do not study it in this work. We keep abstention to avoid re-running all llms and omit rare cases of abstention. Our final RQA prompt is in Prompt A.8.4.
We then designed our QA prompt by mimicking the format of the final RQA prompt, shown in Prompt A.8.4. We initially wrote the constraint that the answer must be “short” and “just a few words,” but we felt these instructions were ambiguous, and the easy and hard entities split of our dataset had answers that were longer than just a few words; as a result, we removed these instructions, and used “the” instead of “a” to make it clear that there is only one valid answer (Prompt A.8.4). After removing these instructions, we noticed that models would often generate very long explanations before or after answering the question. To avoid this, we added an instruction stating that we were just looking for the answer and no explanation (Prompt A.8.4). Our final QA prompt is in Prompt A.8.4.
@@ A.3 Model Details
LLaMA, Mistral, and Yi models are accessed via huggingface, and all other models are accessed through their respective api endpoints. We allocated 8 NVIDIA:A6000s for Mixtral-8x22B, 8 NVIDIA:A5000s for Mixtral-8x7B, Yi-34B, and LLaMA-70B, 2 NVIDIA:A6000s for Yi-9B and LLaMA-8B, and 1 NVIDIA:A6000 for all other non-api models (which were run on CPU only). Each model was allocated 24 hours to run both QA and RQA on our dataset.
@@ A.4 Metric Details
To design a metric for QA accuracy, we consider seven answer equivalence metrics, which check if a candidate answer $a_{cand}$ is semantically equivalent to a ground-truth answer $a_{true}$: 1) DSPy-optimized GPT-4o; 2) A rule-based method designed specifically for each dataset; 3) Exact match; 4) Token F1 score; 5) Token Recall Score; 6) Token Precision Score; and 7) PEDANTS Li et al. (2024b), a classifier designed for answer equivalence. The DSPy method in (1) uses a maximum of 10 bootstrapped demos, a maximum of 10 labeled demos, and 20 candidate programs; it uses 64 examples for training (seeding the prompts) and 64 examples for validation. We decide the optimal decision thresholds for (4), (5), and (6) using the 64 validation examples. We present the agreement with human annotations of each metric in Table 3, which is how we picked the metric to use for each dataset split. In all, our QA accuracy metric has 94% raw agreement with humans on 1152 held-out examples.
Since there are no automated metrics to check whether a question $q$ can correctly be answered by an entity $a$, we design our own metric for RQA accuracy. Given the strength of the DSPy GPT-4o approach in QA accuracy, we similarly design a DSPy-optimized GPT-4o classifier that determines if $q$ is correctly answered by $a$, using the same hyperparameters for QA accuracy. Overall, this RQA accuracy metric has 90% raw agreement with humans on 1152 held-out examples. We also considered Jury approaches Verga et al. (2024), which ensemble multiple llms instead of relying just on a single llm. However, using majority vote with three/five llms boosted our metric’s accuracy by less than 2%, which we did not feel justified the much larger computational expenses.
@@ A.6 RQA with Prompting Engineering
To explore if RQA weaknesses can be alleviated with prompt engineering efforts Schulhoff et al. (2024), we test three prompting strategies: 1) Zero-Shot Chain-of-Thought Prompting (asking the llm to “Think step by step” before answering); 2) Self-Verification (asking the llm to “Check if the question is accurate after generating a question”); and 3) Five-Shot Prompting (including five exemplars showing the model how to generate a question for an answer). To write exemplars for (3), we pick question/answer pairs when RQA succeeds in the zero-shot setting to make the priors in the exemplars most similar to the model’s original generations. The prompts for (1), (2), and (3) are in Prompts A.8.4, A.8.4, and A.8.4, respectively.
We experiment with GPT-4 on Numbers and Numbers+Text, as the model showed a surprising RQA weakness in these settings. GPT-4 is also considered to respond well to prompt engineering efforts, making it a suitable candidate for our prompting strategies. Overall, none of these prompting strategies can close the accuracy gap between RQA and QA (Figure 7). Chain-of-thought prompting increases GPT-4’s RQA accuracy by $\sim 0.15$, but it is still significantly lower than QA, which does not use chain-of-thought. This shows that the accuracy gap between QA and RQA may be an inherent reasoning flaw of current llms that cannot be fully mitigated via prompt engineering.
@@ A.7 Full Consistency Analysis
Figure 8 reports the full consistency analysis for all 16 of our llms and all truth table scenarios. The four categories reported in Figure 3 encompass most of the truth table. Further, even for smaller llms, our claims hold; llms can often detect their own question inaccuracies from RQA through QA.
@@ A.8.1 Are RQA questions unanswerable?
We now seek to understand the types of RQA questions generated in the Number+Text setting, complementing our analysis in section 3.4. The Number+Text questions have higher variance and cannot be as neatly categorized as in section 3.4 (e.g. single-step computation). So instead, we study the answerability of 30 generated questions from each llm, i.e., if the question is clear but leads to an incorrect answer, or if the question has an issue that makes it difficult to answer. We adopt five categories of unanswerable questions from Rogers et al. (2023):
@@
2) No Consensus on the Answer: the question does not have a single, agreed-upon answer . For example, LLaMA generates the question What is the unique property of the Lie algebra E8 that makes it particularly interesting in theoretical physics?, but Lie algebra has many distinct, interesting properties that would answer the question.
3) Information not yet Discovered: the answer to this question is not yet known. For example, GPT-4 generates the question How long, in terms of word count, is the sentence that holds the record for being the longest in the English language without using any punctuation?, but it is not yet known what could theoretically be the longest sentence.
4) Missing Information: the question does not have enough information, or it is too vague. For example, GPT-4 generates the question How many individuals attended the annual community festival last year according to the final headcount?, which cannot be answered without knowing more details.
@@
As expected, when RQA succeeds, questions are mostly answerable (Figure 9). However, a non-trivial proportion of generated questions when RQA fails are unanswerable, reaching nearly 60% for GPT-4. The most common types of unanswerable questions are those that are missing information, meaning that they are too vague or ambiguous, or those that have false premises or assumptions. While several works explore methods to answer ambiguous questions Min et al. (2020); Kim et al. (2023a) or questions with false presuppositions Yu et al. (2023); Kim et al. (2023b), our analysis reveals a need to avoid generating ambiguous or faulty-presupposition questions in RQA.
@@ A.8.2 Do llms give the same RQA questions?
Interestingly, models in the same family are more likely to generate duplicated questions. For example, GPT-3.5, GPT-4, and GPT-4o generate the same questions in RQA more often than when compared to other llm families. Thus, we speculate that these model families likely share similar pre-training and alignment data, which is optimized on through different training recipes.
@@ A.8.4 Are any RQA questions memorized?
Since the duplicates in Appendix A.8.3 suggest that llms may just be retrieving similar questions from pretraining rather than reasoning towards new questions in RQA, we now investigate the novelty of the RQA questions Merrill et al. (2024), i.e., whether they are exactly copied from pretraining. We do not know which corpora all of our llms are trained on, so we use the Dolma Soldaini et al. (2024) corpus as a proxy for pretraining data. For each generated RQA question $\hat{q}$, we compute how frequently the exact question $\hat{q}$ appears in Dolma via infini-gram Liu et al. (2024b).
Table 7 reveals in total, $2.87$% of RQA questions are exactly found in Dolma. For comparison, $1.25$% of our ground-truth questions exist in Dolma. While we did not explicitly prompt the model to give a new question that it has not seen in pretraining, practitioners may need to design specialized techniques if they desire novel RQA questions.
When comparing exact question match frequency by model, weaker/smaller llms tend to copy more from pretraining data, suggesting that smaller llms are more prone to RQA memorization. Further, the Hard Fact setting is much less prone to question copying in RQA, likely because the RQA input answers have very low pretraining token count (section 3.3), which further supports that llms may struggle to retrieve exact pretraining knowledge for long-tail facts Kandpal et al. (2023).